Solve and check each of the following equations for x.
(2)/(x+ 2)=(3)/(x+ 6)+(9)/(x^2 +8x+ 12)

Answer by scott8148(5892)

x^2+8x+12 is LCD, so x cannot equal -2 or -6 (would result in zero denominator)

2(x+6)=3(x+2)+9 ... 2x+12=3x+15